### v2

# On the stability of the penalty function for the $\mathbb{Z}^{2}$-hard square shift

Abstract.
We investigate the stability of maximizing measures for a penalty function of a two-dimensional subshift of finite type, building on the work of Gonschorowski et al. [GQS]. In the one-dimensional case, such measures remain stable under Lipschitz perturbations for any subshift of finite type. However, instability arises for a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic point and zero entropy. This raises the question of whether stability persists in two-dimensional subshifts of finite type with positive topological entropy. In this paper, we address this question by studying the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift with positive entropy. Our main theorem establishes that, in contrast to previous results, a penalty function of the hard square shift remains stable under Lipschitz perturbations.
Key words and phrases:
2010 Mathematics Subject Classification:
Primary 37B51, 37B10, 37B25
1. Introduction
Ergodic optimization is the study of maximizing measures. In its most basic form, let $T:X\rightarrow X$ be a continuous map on a compact metric space $X$ and for a continuous function $\varphi:X\rightarrow\mathbb{R}$ we consider the maximum ergodic average
[…]
where $\mathcal{M}_{T}(X)$ is the space of $T$-invariant Borel probability measures on $X$ endowed with the weak*-topology. An invariant measure which attains the maximum is called a maximizing measure for $\varphi$ and denote by $\mathcal{M}_{{\rm max}}(\varphi)$ the set of maximizing measures for $\varphi$.
The stability of maximizing measures for a penalty function of a subshift of finite type was established by Gonschorowski et al. [GQS]. A penalty function is defined on the forbidden set of a subshift of finite type, assigning a value of 00 to admissible local configurations near the origin and $-1$ otherwise (see §2 for more details). It is straightforward to see that every maximizing measure of a penalty function is supported on the given subshift of finite type. In the one-dimensional case, maximizing measures remain supported on the given subshift under Lipschitz perturbations for any subshift of finite type. However, in the two-dimensional case, there exists a subshift of finite type where this stability fails.
In [GQS], the authors highlight the difference between one and two dimensions, demonstrating that instability arises in a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic points and zero entropy. This naturally prompts the question of whether stability persists for a two-dimensional subshift of finite type with positive topological entropy. In this paper, we address this question by investigating the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift of finite type with positive entropy (See [Pav12] for more properties of hard square shifts). Our main theorem establishes that, in contrast to the results on $\mathbb{Z}^{2}$ subshift of finite type as presented by Gonschorowski et al., the penalty function of the hard square shift remains stable under Lipschitz perturbations.
Informally, a subshift of finite type is defined by specifying a finite set of finite “forbidden patterns” $F$ made up of letters from alphabet $\mathcal{A}$, and defining $X_{F}$ to be the set of configurations in $\mathcal{A}^{\mathbb{Z}^{d}}$ in which no pattern from $F$ appears (see §2 for more details). The set $F$ is called a forbidden set. The $\mathbb{Z}^{2}$-hard square shift is a subshift of finite type where no two adjacent 1’s appear, either horizontally or vertically. Although other forbidden sets can be used to define the hard square shift, we will use the following forbidden set $F$ to define the penalty function later:
(1) $F=\left\{\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\right\}.$
We now define the penalty function as follows:
$f(x)=\left\{\begin{array}[]{cc}-1&\mbox{if}~{}\begin{smallmatrix}x_{(0,1)}&x_{(1,1)}\\ x_{(0,0)}&x_{(1,0)}\end{smallmatrix}\in F\\ 0&\mbox{otherwise}\end{array}\right..$
Now we can state our main theorem.
Theorem 1.
Let $X$ be the hard square shift and $f$ be the penalty function. Then there exists $\varepsilon>0$ such that for every Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$, every maximizing measure of $g$ is supported on $X$.
We remark that the stability result for the hard square shift is relatively straightforward, since forbidden words can be easily eliminated by replacing $1$ with 00. However, extending this result to other nearest neighbor subshifts of finite type, such as those discussed in [Pav12], appears to be more challenging. This difficulty arises because, in general, there is no method to identify the exact locations of forbidden words, making it impossible to estimate the distances required to remove them.
For the remainder of this paper, we fix our notations and definitions in §2 and provide the proof of the main theorem in §3.
2. Settings
Let $\mathcal{A}$ be a finite set, which we call an alphabet. The $\mathbb{Z}^{2}$ full shift on $\mathcal{A}$ is the set $\mathcal{A}^{\mathbb{Z}^{2}}$, endowed with the product topology of the discrete topology. Define a metric by
$d(\underline{x},\underline{y})=\left\{\begin{array}[]{cc}\frac{1}{2^{i}}&\underline{x}\neq\underline{y}\\ 0&\mbox{otherwise}\end{array}\right.$
for $\underline{x},\underline{y}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where $i=\inf\{\|u\|_{\infty}:x_{u}\neq y_{u}\}$. Then, this metric is compatible with the product topology.
For any full shift $\mathcal{A}^{\mathbb{Z}^{2}}$, we define the $\mathbb{Z}^{2}$-action $\{\sigma_{u}\}_{u\in\mathbb{Z}^{2}}$ on $\mathcal{A}^{\mathbb{Z}^{2}}$ as follows: for any $u\in\mathbb{Z}^{2}$ and $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$, $(\sigma_{u}(\underline{x}))_{v}=x_{u+v}$ for all $u\in\mathbb{Z}^{2}$.
A subset $X\subset\mathcal{A}^{\mathbb{Z}^{2}}$ is a subshift if it is closed and shift-invariant, i.e. for any $\underline{x}$ and $u\in\mathbb{Z}^{2}$, $\sigma_{u}(\underline{x})\in X$.
A configuration $w$ on the alphabet $\mathcal{A}$ is any mapping from a non-empty subset $S$ of $\mathbb{Z}^{2}$ to $\mathcal{A}$, where $S$ is called the shape of $w$. For any configuration $w$ with shape $S$ and any $T\subset S$, denote by $w|_{T}$ the restriction of $w$ to $T$, i.e. the subconfiguration of $w$ appearing $T$.
A subshift $X$ is a shift of finite type if there exists a finite collection $F$ of finite configuration on $\mathcal{A}$ such that $X=X_{F}$ to be the set of $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$ such that $\underline{x}|_{S}\notin F$ for all finite $S\subset\mathbb{Z}^{2}$.
Definition 2.1 (The $\mathbb{Z}^{2}$-hard square shift).
Let $\mathcal{A}=\{0,1\}$ and set the forbidden set by
$F=\left\{\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\right\}.$
Then the subshift $X=X_{F}$ is called the $\mathbb{Z}^{2}$-hard square shift.
It is easy to see that the topological entropy of the hard square shift is positive.
For any $a,b\in\mathbb{Z}$ with $a<b$, we use $[a,b]$ to denote $\{a,a+1,\ldots,b\}$. For each $n\geq 0$ define the box of size $n$ as
$\Lambda_{n}=[-n,n]\times[-n,n].$
The cardinality of $\Lambda_{n}$ is given by $\lambda_{n}=\#\Lambda_{n}=(2n+1)^{2}$. For a continuous function $f$ and a nonempty subset $T\subset\mathbb{Z}^{2}$ define a dynamical sum over $T$ by
$S_{T}f=\sum_{u\in T}f\circ\sigma^{u}.$
3. Proof of the main theorem
First we recall the following Lemma.
Lemma 3.1 ([GQS, Lemma 2.1.]).
Let $J\subset X$ be a subset of a compact metric space $X$ and $f$ be a Lipschitz continuous function with $f|_{J}=0$. For $\varepsilon>0$ and a Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$ we have
$|g(x)-g(y)|<\varepsilon d(x,y)$
for all $x,y\in J$.
This lemma will be applied in our setting with $B=f^{-1}\{0\}$. With this preparation, we now proceed to the proof of our main theorem. A key feature of this proof is its extension of the coupling and splicing argument, as well as the ”path-wise surgery” technique from [GQS], to a two-dimensional case.
Proof of Theorem 1.
Let $\varepsilon=\frac{1}{64}$ and $g$ be a Lipschitz function with $\|f-g\|_{{\rm Lip}}<\varepsilon$.
Set $I=f^{-1}\{0\}$. Since the set of maximizing measures is convex and closed, it suffices to prove the result for ergodic measures. Let $\mu$ be an ergodic invariant measure supported on $X^{c}$.
(Case 1). $\mu(I^{c})\geq 1/2$.
[…]
which completes the proof.
(Case 2). $\mu(I^{c})\leq 1/2$.
Let $\underline{x}$ be a generic point for the measure $\mu$. For each $i\in\mathbb{Z}$ let
$S^{i}=\{u_{1}\in\mathbb{Z}:\sigma^{(u_{1},i)}\underline{x}\in I^{c}\}.$
For $u_{1},v_{1}\in S^{i}$ set a relation $u_{1}rv_{1}$ by
$u_{1}rv_{1}\Leftrightarrow|u_{1}-v_{1}|=1.$
Moreover define the equivalent relation $\sim$ on $S^{i}$ by
$u_{1}\sim v_{1}\Leftrightarrow\ \mbox{there exist}\ w_{1}^{1},\ldots,w_{1}^{p}\in S^{i}\ \mbox{such that}\ u_{1}rw_{1}^{1};w_{1}^{1}rw_{1}^{2};\cdots;w_{1}^{p}rv_{1}.$
Then we get the sequence of bad words on ×$\mathbb{Z}\times\{i\}$. Set $S^{i}/\sim=\{B_{n}^{i}\}_{n\in\mathbb{Z}}$ where $B_{0}^{i}$ is the equivalent class including 00 if $0\in S^{i}$and it is the equivalent class including $\min\{u_{1}>0:(u_{1},i)\in S^{i}\}$ if $0\notin S$. Setting $\alpha^{i}_{n}=\min\{u_{1}\in B_{n}^{i}\}$ and $\beta_{n}^{i}=\max\{u_{1}\in B_{n}^{i}\}$ for each $n$, we have $B_{n}^{i}=[\alpha^{i}_{n},\beta^{i}_{n}]$.
Then we define a sequence of configurations by inductively replacing $1$ to 00 on each line. First set, $\underline{x}^{(-1)}:=\underline{x}$ and define $\underline{x}^{(0)}$ by
$x^{(0)}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{0}\ \mbox{and}\ u_{2}=0\\ x_{(u_{1},u_{2})}&\mbox{othewise}.\end{array}\right.$
For $k\geq 1$, we define $\underline{x}^{(2k-1)}$ and $\underline{x}^{(2k)}$ inductively as follows:
$x^{(2k-1)}_{(u_{1},u_{2})}$ =$=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k}\ \mbox{and}\ u_{2}=k\\ x^{2k}_{(u_{1},u_{2})}&\mbox{otherwise},\end{array}\right.$
$x^{(2k)}_{(u_{1},u_{2})}$ =$=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{-k}\ \mbox{and}\ u_{2}=-k\\ x^{2k-1}_{(u_{1},u_{2})}&\mbox{otherwise}.\end{array}\right.$
Then there is no bad word in $\underline{x}^{(2k-1)}$ on $[-k+1,k]\times\mathbb{Z}$ and in $\underline{x}^{(2k)}$ on $[-k,k]\times\mathbb{Z}$.
Furthermore, define $\tilde{x}$ by
$\tilde{x}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k},u_{2}=k\ \mbox{and}\ k\in\mathbb{Z}\\ x_{(u_{1},u_{2})}&\mbox{otherwise}.\end{array}\right.$
Clearly, $\tilde{x}\in X$. By definition, we have $\lim_{k\to\infty}\underline{x}^{(k)}=\tilde{x}$.
Fix sufficiently large $N\geq 1$. We now consider the difference between the dynamical sums of $\underline{x}$ and $\tilde{\underline{x}}$ over $\Lambda_{N}$ :
(2) $S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=\sum_{i=0}^{2N-1}\left(S_{\Lambda_{N}}g(\underline{x}^{i-1})-S_{\Lambda_{N}}g(\underline{x}^{i})\right)+S_{\Lambda_{N}}g(\underline{x}^{(2N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}}).$
The last two terms can be bounded as follows:
(3) $S_{\Lambda_{N}}g(\underline{x}^{(2N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=2\sum_{i=1}^{N}\frac{(2N+1)}{2^{i}}\leq 2(2N+1).$
In order to provide an upper bound for the summation term, we analyze the difference between the dynamical sums of $\underline{x}^{i-1}$ and $\underline{x}^{i}$ over $\Lambda_{N}$ by considering contributions from bad words and good words separately.
Estimate on bad words: Let $i$ be a nonnegative integer. For $n\in\mathbb{Z}$ and $u_{1}\in B_{n}^{i}$ we have $g(\sigma^{(u_{1},i)}\underline{x}^{(i-1)})<-1+\varepsilon$ and $g(\sigma^{(u_{1},i)}\underline{x}^{(i)})>-\varepsilon$. Hence we have
$S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{i-1})-S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{i})<|B_{n}^{i}|(-1+\varepsilon)+|B_{n}^{i}|\varepsilon=|B_{n}^{i}|(-1+2\varepsilon)$
where $|E|$ denotes the cardinality of $E$.
Estimate on Good words: To establish an estimate for good words, we introduce the following notations: For a nonnegative integer $i$ set $I^{i}_{N}$ by
$I^{i}_{N}=\{n:B_{n}^{i}\cap[-N,N]\}$
and let $\ell_{N}^{i}=\min I_{n}^{i}$ and $r_{N}^{i}=\max I_{n}^{i}$. Since we are interested only in bad words within $\Lambda_{N}$, we modify the endpoints as follows. For $n\in[\ell_{N}^{i}-1,r_{N}^{i}+1]$ set
$\tilde{\alpha}_{n}^{i}=\left\{\begin{array}[]{ccc}-N&\mbox{if}&-N\in S^{i}\ \mbox{and}\ n=\ell_{N}^{i}\\ N+1&\mbox{if}&n=r_{N}^{i}+1\\ \alpha_{n}^{i}&\mbox{else}&\end{array}\right.$
and
$\tilde{\beta}_{n}^{i}=\left\{\begin{array}[]{ccc}-N&\mbox{if}&n=\ell_{N}^{i}-1\\ N&\mbox{if}&N\in S^{i}\ \mbox{and}\ n=r_{N}^{i}\\ N+1&\mbox{if}&N\notin S^{i}\ \mbox{and}\ n=r_{N}^{i}+1\\ \beta_{n}^{i}&\mbox{else.}\end{array}\right.$
Refer to Figure 1 for a visual representation of this operation.
[…]
Case1111-i : $-N\in S^{i}$
[…]
Case1111-ii : $-N\notin S^{i}$
[…]
Case2222-i : $N\in S^{i}$
[…]
Case2222-ii : $N\notin S^{i}$
Figure 1. The value of $\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}$
For evaluating the difference between the dynamical sums of $\underline{x}^{i-1}$ and $\underline{x}^{i}$ over “good words”, we use an upper bound on the distance between $\sigma^{u}\underline{x}^{i-1}$ and $\sigma^{u}\underline{x}^{i}$ for each $u\in\Lambda_{N}\setminus\left(\bigcup_{n\in I^{i}_{N}}B_{n}^{i}\right)$. This upper bound depends on the distance to bad words, and this modification does not make that distance any smaller. Let $c^{i}_{n}=\lfloor\frac{\tilde{\alpha}_{n}^{i}-\tilde{\beta}_{n-1}^{i}}{2}\rfloor(\geq 0)$. Define the sets
$G_{n}^{i,+}$ $=([\tilde{\beta}_{n-1}^{i},\tilde{\beta}_{n}^{i}]\times[i,N])\setminus([\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}]\times\{i\})$
and
$G_{n}^{i,-}$ $=[\tilde{\beta}_{n-1}^{i},\tilde{\beta}_{n}^{i}]\times[-N,i-1].$
Then we obtain
$\Lambda_{N}=\bigcup_{n=\ell_{N}^{i}}^{r_{N}^{i}+1}(G_{n}^{i,+}\cup G_{n}^{i,-})\cup\bigcup_{n=\ell_{N}^{i}}^{r_{N}^{i}}[\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}].$
Now, let $0\leq i\leq N$ and consider $n\in[l_{N}^{-i},r_{N}^{-i}]$ satisfying $0\leq c_{n}^{-i}<N+i$. For $(u_{1},u_{2})\in G_{n}^{-i,+}$, the distance
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})$
is determined by three cases, and the computation is divided into four regions:
$A$ =$=\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+1\leq u_{1}\leq\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i},-i\leq u_{2}\leq-i+u_{1}-(\tilde{\beta}_{n-1}^{-i}+1)\}$
$\quad\cup\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i}+1\leq u_{1}\leq\tilde{\alpha}_{n}^{-i}-1,-i\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1-u_{1}+(\tilde{\beta}_{n-1}^{-i}+\tilde{c}_{n}^{-i}+1)\};$
$B$ =$=\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+1\leq u_{1}\leq\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i},-i+u_{1}-(\tilde{\beta}_{n-1}^{-i}+1)\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1\}$
$\quad\cup\{(u_{1},u_{2}):\tilde{\beta}^{-i}_{n-1}+\tilde{c}_{n}^{-i}+1\leq u_{1}\leq\tilde{\alpha}_{n}^{-i}-1,-i+u_{1}-(\tilde{\alpha}_{n}^{-i}-1)\leq u_{2}\leq-i+\tilde{c}_{n}^{-i}-1\};$
$C$ =$=\{(u_{1},u_{2}):\tilde{\alpha}_{n}^{-i}\leq u_{1}\leq\tilde{\beta}_{n}^{-i},-i+1\leq u_{2}\leq\tilde{c}_{n}^{-i}\};$
$D$ $=[\tilde{\beta}_{n-1}^{-i}+1,\tilde{\beta}_{n}^{-i}-1]\times[-i+\tilde{c}_{n}^{-i}+1,N].$
To illustrate this, we assign letters to each area as shown in Figure 2. The red graph represents a path where both $u_{1}$ and $u_{2}$ increase by 1 at each step.
For $(u_{1},u_{2})$ in the region $A$, the horizontal distance from the bad words is the determining factor. Specifically,
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})=\left\{\begin{array}[]{ccc}\frac{1}{2^{u_{1}-\tilde{\beta}_{n-1}^{-i}}}&\mbox{if}&u_{1}\leq\tilde{\beta}_{n-1}^{-i}+c_{n}^{-i}\\ \frac{1}{2^{\tilde{\alpha}_{n}^{-i}-u_{1}}}&\mbox{if}&u_{1}>\tilde{\beta}_{n-1}^{-i}+c_{n}^{-i}.\end{array}\right.$
For $(u_{1},u_{2})$ in the regions $B,C$ and $D$, the vertical distance is the determining factor. Specifically,
$d(\sigma^{(u_{1},u_{2})}\underline{x}^{(2i-1)},\sigma^{(u_{1},u_{2})}\underline{x}^{(2i)})=\frac{1}{2^{u_{2}}}.$
Taking into account the symmetry of regions $A$ and $B$, we compute as follows. By Lemma 3.1, we obtain the following bound:
$S_{G_{n}^{-i,+}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,+}}g(\underline{x}^{(2i)})$ $<\left(\sum_{k=1}^{c_{n}^{i}}2k\cdot\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}2k\cdot\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\frac{1}{2^{k}}\right.$
$\hskip 30.00005pt\left.+\sum_{k=c_{n}^{i}+1}^{N-i}(\tilde{\beta}^{i}_{n}-\tilde{\beta}^{i}_{n-1})\frac{1}{2^{k}}+\sum_{k=1}^{c_{n}^{i}}\frac{2}{2^{k}}\right)\varepsilon$
$<\left(\sum_{k=1}^{c_{n}^{i}}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}+\sum_{k=c_{n}^{i}+1}^{N-i}\frac{(\tilde{\alpha}^{i}_{n}-\tilde{\beta}_{n-1})}{2^{k}}\right)\varepsilon$
$\hskip 44.0pt(\because\tilde{\beta}_{n}^{i}-\tilde{\beta}_{n-1}^{i}=\tilde{\beta}_{n}^{i}-\tilde{\alpha}_{n}^{i}+\tilde{\alpha}_{n}^{i}-\tilde{\beta}_{n-1}^{i})$
$\leq\left(\sum_{k=1}^{c_{n}^{i}}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}+\sum_{k=c_{n}^{i}+1}^{N-i}\frac{2k}{2^{k}}\right)\varepsilon$
$\hskip 44.0pt(\because\tilde{\alpha}^{i}_{n}-\tilde{\beta}_{n-1}\leq 2c_{n}^{i}+1)$
$\leq\left(\sum_{k=1}^{N-i}\frac{4k+2}{2^{k}}+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\sum_{k=1}^{N-i}\frac{1}{2^{k}}\right)\varepsilon$
(4) $=\left(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})\right)\varepsilon.$
$-i$$\tilde{\alpha}_{n}^{-i}$$\tilde{\beta}_{n}^{-i}$$\tilde{\beta}_{n-1}^{-i}$$c_{n}^{-i}$$\tilde{\beta}_{n-1}^{-i}+c_{n}^{-i}$
$\approx$
[…]
Figure 2. The value of point in good block
For $n\in[\ell_{N}^{-i},r_{N}^{-i}]$ such that $N\leq c_{n}^{-i}\leq N+i$ there are no regions D and A in Figure2, and region B is cut off in the middle. Hence it is easy to see that we have
$S_{G_{n}^{-i,-}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,-}}g(\underline{x}^{(2i)})$ $<(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n}))\varepsilon.$
By the same argument for $10\leq i\leq N-$ 1 and $n\in[r_{N}^{i},\ell_{N}^{i}]$ we have
(5) $S_{G_{n}^{i,\pm}}g(\underline{x}^{2i})-S_{G_{n}^{i,\pm}}g(\underline{x}^{2i+1})$ $<(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n})).$
Noting that $G_{n}^{-i,+}\cap G_{n}^{-i,+}=\emptyset$, we have
$\sum_{i=0}^{2N}\left(S_{\Lambda_{N}}g(\underline{x}^{i-1})-S_{\Lambda_{N}}g(\underline{x}^{i})\right)$ $\leq\sum_{i=0}^{N}\sum_{n=\ell_{N}^{-i}}^{r_{N}^{-i}}\left(S_{G_{n}^{-i,+}\cup G_{n}^{-i,-}}g(\underline{x}^{(2i-1)})-S_{G_{n}^{-i,+}\cup G_{n}^{-i,-}}g(\underline{x}^{(2i)})\right)$
$\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\left(S_{G_{n}^{i,+}\cup G_{n}^{i,-}}g(\underline{x}^{(2i)})-S_{G_{n}^{i,+}\cup G_{n}^{i,-}}g(\underline{x}^{(2i+1)})\right)$
$\quad\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{-i}}^{r_{N}^{-i}}\left(S_{B_{n}^{-i}\times\{-i\}}g(\underline{x}^{(2i-1)})-S_{B_{n}^{-i}\times\{-i\}}g(\underline{x}^{(2i)})\right)$
$\quad\quad\quad+\sum_{i=0}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\left(S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{(2i)})-S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{(2i+1)})\right)$
$<2\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}(14+(\tilde{\beta}^{i}_{n}-\tilde{\alpha}^{i}_{n}))\varepsilon+\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}(-1+2\varepsilon)$
$\leq 28\varepsilon\sum_{i=-N}^{N}\#I_{N}^{i}+2\varepsilon\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}+\sum_{i=-N}^{N}\sum_{n=\ell_{N}^{i}}^{r_{N}^{i}}\#B_{n}^{i}(-1+2\varepsilon)$
$\leq(28\varepsilon+2\varepsilon-1+2\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}$
$=(-1+32\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}.$
Dividing by $(2N+1)^{2}$ by the both sides of (2), we have
$\frac{1}{(2N+1)^{2}}\left(S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}})\right)$ $<\frac{2}{(2N+1)}+\frac{1}{(2N+1)^{2}}(-1+32\varepsilon)\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}$
$\leq\frac{2}{(2N+1)}-\frac{1}{2}\frac{1}{(2N+1)^{2}}\#\{u\in\Lambda_{N}:\sigma^{u}\underline{x}\in I^{c}\}.$
Hence we have
$\liminf_{N\to\infty}\frac{1}{(2N+1)^{2}}S_{\Lambda_{N}}g(\underline{x})+\frac{1}{2}\mu(I^{c})$ $<\liminf_{N\to\infty}\frac{1}{(2N+1)^{2}}S_{\Lambda_{N}}g(\tilde{\underline{x}}).$
Since $\underline{x}$ is a generic point of $\mu$ and we see that there exists an invariant probability measure $\nu$ with support in $X$ by passing to a subsequence of the sequence of empirical measures for $\tilde{x}$, we have
[…]
, which complete the proof.
Acknowledgements.
The second author was partially supported by JSPS KAKENHI Grant Number 21K13816.
Data Availability.  Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.